# Cutting a Pancake with an Exotic Knife

### v1

@@ 9.1 The constrained long-legged $\overline{{\sf T}}$
To prove that this formula is correct, ideally we would like an explicit construction that achieves it for all $n\geq 0$, and a proof that no greater value is possible. At present we have neither of these things for $n>4$, so we leave (21) as a conjecture.
Figure 19: Two constrained $\overline{\sf A}$’s can intersect in 8 points, and divide the plane into $a_{\overline{\sf A}}(2)=13$ regions. (In this picture one $\overline{\sf A}$ has a purple tip, the other a brown tip. The arrowheads are not shown.)
@@ 10.1 Convex polygons
The solutions for $k=3$, 44, and $5$ (for the triangle, square or convex quadrilateral, and pentagon) are given in sequences A077588, A069894, and A386485. It was a slight surprise to discover that taking $k=1$ and $k=2$ in (24) matched A386480 and A051890, which are the numbers of regions attainable with $n$ circles (see §10.3) and $n$ ellipses, respectively. Most surprising, however, was the case $k=8$, which we discuss in the next subsection.
@@ 10.2 The octagon and concave quadrilateral
Figure 23: Two concave quadrilaterals also divide the plane into 18 regions. Is it just a concidence tht the maximum number of pieces that can be obtained with $n$ (convex) octagonal cuts is the same as can be obtained with $n$ concave quadrilateral cuts? We do not know!
@@ 10.5 The pentagram and hexagram
.A regular pentagram or hexagram is based on 5 or 6 equally spaced points around a circle. To attain the maximum number of regions we will use regular figures, but our analysis applies equally to non-regular figures. A single pentagram (regular or non-regular) has 5 base nodes, the exterior vertices, with degree $d_{B}=2$, and 5 self-intersections, in the interior. Two pentagrams can meet in at most 20 points, so with $n$ pentagrams we have $V_{{\mathcal{C}}}\leq 5n+20\binom{n}{2}$. For $n$ pentagrams our basic equation (10) then gives $R\leq V_{{\mathcal{C}}}+2\leq 5n+20\binom{n}{2}+2.=10n^{2}-5n+2$ (A383466), a bound which is easily achieved using regular pentagons based on $5n$ equally spaced points around a circle. The construction is illustrated for $n=3$ in Fig. 26. To verify the number of regions, note that each of the base nodes is a vertex of a triangular region (indicated by the three blue dots in Fig. 26) containing 2n12n-1 cells. We were led to investigate pentagrams because an optimal 5-chain (§5) necessarily contains a pentagram (see Figs. 7 and 10).
A hexagram initially contains two disjoint triangles and is not a connected graph, which would violate the assumption that our shapes are connected, but is connected if we take the self-intersections into account. The analysis is similar to that of the pentagram, so we just state the result: the maximum number of regions that can be attained in the plane by drawing $n$ hexagrams is $2(6n^{2}-3n+1)$ (A386477).
@@ 11 The lollipop
This can be achieved for $n\leq 3$ (Fig. 28), though it is already unclear what is happening when $n\geq 4$. As in §9.2, computational methods may be used, but the delicate nature of the optimal solutions makes them highly sensitive to numerical errors.
@@ 12 Open problems
    More generally, for any shape $S$ for which we have determined the maximum number of regions $a_{S}(n)$ (which includes all the entries in Tables 1 and 2), we may ask for the number of distinct graphs having that many regions. The case when $S$ is a circle has been studied by J. Wild in A250001 and A288554. He shows that the numbers of distinct planar graphs composed of $n$ overlapping circles that divide the plane into $n^{2}-n+2$ regions are respectively $1,1,4,45,5102$ for $n=1,\ldots,5$. The four possibilities for $n=3$ are shown in Fig. 29. One of the $45$ is shown in Fig. 24).

### v2

@@ 9.1 The constrained long-legged $\overline{{\sf T}}$
To prove that this formula is correct, ideally we would like an explicit construction that achieves it for all $n\geq 0$, and a proof that no greater value is possible. At present we have neither of these things for $n>4$, so we leave (22) as a conjecture. Numerical coordinates for the best configurations we have found for $n\leq 12$ can be found in A389614.
Figure 23: Two constrained $\overline{\sf A}$’s can intersect in 8 points, and divide the plane into $a_{\overline{\sf A}}(2)=13$ regions. (In this picture one $\overline{\sf A}$ has a purple tip, the other a brown tip. One of the intersection points is not shown: it occurs a long way below the bottom of the picture. The arrowheads are also not shown.)
@@ 10.1 Convex polygons
The solutions for $k=3$, 44, and $5$ (for the triangle, square or convex quadrilateral, and pentagon) are given in sequences A077588, A069894, and A386485. It was a slight surprise to discover that taking $k=1$ and $k=2$ in (25) matched A386480 and A051890, which are the numbers of regions attainable with $n$ circles (see §10.3) and $n$ ellipses, respectively. More surprising, however, was the case $k=8$, which we discuss in the next subsection.
@@ 10.2 The octagon and concave quadrilateral
Figure 27: Two concave quadrilaterals also divide the plane into 18 regions. Is it just a concidence tht the maximum number of pieces that can be obtained with $n$ (convex) octagonal cuts is the same as what can be obtained with $n$ concave quadrilateral cuts? We do not know!
@@ 10.5 The pentagram and hexagram
.A regular pentagram or hexagram is based on 5 or 6 equally spaced points around a circle. To attain the maximum number of regions we will use regular figures, but our analysis applies equally to non-regular figures. A single pentagram (regular or non-regular) has 5 base nodes, the exterior vertices, with degree $d_{B}=2$, and 5 self-intersections, in the interior. Two pentagrams can meet in at most 20 points, so with $n$ pentagrams we have $V_{{\mathcal{C}}}\leq 5n+20\binom{n}{2}$. For $n$ pentagrams our basic equation (10) then gives $R\leq V_{{\mathcal{C}}}+2\leq 5n+20\binom{n}{2}+2.=10n^{2}-5n+2$ (A383466), a bound which is easily achieved using regular pentagons based on $5n$ equally spaced points around a circle. The construction is illustrated for $n=3$ in Fig. 30. To verify the number of regions, note that each of the base nodes is a vertex of a triangular region (indicated by the three blue dots in Fig. 30), containing 2n12n-1 cells. We were led to investigate pentagrams because an optimal 5-chain (§5) necessarily contains a pentagram (see Figs. 7 and 10).
A hexagram initially contains two disjoint triangles and is not a connected graph, which would violate the assumption that our shapes are connected, but it is connected if we take the self-intersections into account. The analysis is similar to that of the pentagram, so we just state the result: the maximum number of regions that can be attained in the plane by drawing $n$ hexagrams is $2(6n^{2}-3n+1)$ (A386477).
@@ 11 The lollipop
This can be achieved for $n\leq 3$ (Fig. 32), but even for $n=4$ we are unable to get close to the upper bound of 47. As in §9.2, computational methods may be used, but the delicate nature of the optimal solutions makes them highly sensitive to numerical errors.
@@ 12 Open problems
    More generally, for any shape $S$ for which we have determined the maximum number of regions $a_{S}(n)$ (which includes all the entries in Tables 1 and 2), we may ask for the number of distinct graphs having that many regions. The case when $S$ is a circle has been studied by J. Wild in A250001 and A288554. Wild shows that the numbers of distinct planar graphs composed of $n$ overlapping circles that divide the plane into $n^{2}-n+2$ regions are respectively $1,1,4,45,5102$ for $n=1,\ldots,5$. The four possibilities for $n=3$ are shown in Fig. 33. One of the $45$ is shown in Fig. 28. Arrangements of circles in the plane are classified from a different point of view in Mathar [11].